# HateSieve: A Contrastive Learning Framework for Detecting and Segmenting Hateful Content in Multimodal Memes

### v1

@@ Abstract
Amidst the rise of Large Multimodal Models (LMMs) and their widespread application in generating and interpreting complex content, the risk of propagating biased and harmful memes remains significant. Current safety measures often fail to detect subtly integrated hateful content within “Confounder Memes”. To address this, we introduce HateSieve, a new framework designed to enhance the detection and segmentation of hateful elements in memes. HateSieve features a novel Contrastive Meme Generator that creates semantically paired memes, a customized triplet dataset for contrastive learning, and an Image-Text Alignment module that produces context-aware embeddings for accurate meme segmentation. Empirical experiments on the Hateful Meme Dataset show that HateSieve not only surpasses existing LMMs in performance with fewer trainable parameters but also offers a robust mechanism for precisely identifying and isolating hateful content. Caution: Contains academic discussions of hate speech; viewer discretion advised.
@@ 1 Introduction
The emergence of Large Multimodal Models (LMMs), such as GPT-4V Achiam et al. (2023), Stable Diffusion Rombach et al. (2022), and DALL·E Ramesh et al. (2022), has ushered in a new era where people increasingly rely on these models for generating and interpreting visual and textual information. While these services simplify access to information, they also introduce the risk of unregulated content that could distort public perception and cause harm to social groups Su et al. (2023); Qu et al. (2023); Chin et al. (2023); Qu et al. (2024); Meng Huat Tiong et al. (2024); Lin et al. (2024). To address this risk, current LMM platforms implement safety filters, incorporating Alignment Ghafouri et al. (2023), Inference Guidance Chiang et al. (2023), and Input&Output Filter Alon and Kamfonas (2023) to detect and eliminate offensive or inappropriate components in both images and text.
However, these safety filters face challenges in identifying “Confounder Memes” Kiela et al. (2020); Mei et al. (2023), which deliberately combine visual and textual elements to convey biased and discriminatory messages. Such memes may not have obvious offensive content in their individual components, instead delivering their harmful messages through their combined presentation.
@@
Figure 2: An overview of our HateSieve Framework
A straightforward solution involves supervised fine-tuning (SFT) of LMMs Lin et al. (2024) to recognize hateful semantics in memes. Nevertheless, this approach encounters obstacles: (1) the scarcity of detailed annotations in existing hateful meme datasets, making it difficult for models to accurately distinguish between hateful and non-hateful memes, especially when the differences are subtle. (2) Deploying LMMs as safety filters alongside their regular online service111”Online service” refers to real-time applications like chatbots, virtual assistants, and image recognition platforms that use LMMs. usage is computationally intensive and non-trival Lin et al. (2024). Alternatively, a lightweight classifier Kumar and Nandakumar (2022a); Mei et al. (2023) could be trained from scratch using a specialized hateful meme dataset, but this method suffers from limited interpretability and cannot provide a clear rationale for its classifications.
To address these challenges, we introduce HateSieve, a new framework for detecting hateful memes, detailed in Figure 2. HateSieve features an Image-Text Alignment (ITA) module that generates context-aware embeddings using contrastive learning to differentiate between semantically similar memes with varying sentiments. Owing to the absence of suitable annotations in existing datasets, we developed a contrastive meme generator that produces related triplet pairs for training. Pre-training the ITA on this dataset enables it to segment hateful content across image and text modalities effectively, with subsequent fine-tuning on classification tasks demonstrating enhanced performance over current models.
  1. 1.
    We propose a unique contrastive meme generator that synthesizes semantically similar triplet pairs, addressing the gap where specific pairwise annotations are lacking in existing datasets.
  2. 2.
    Our ITA module efficiently produces context-aware embeddings for images and texts, improving the detection and segmentation of hateful content.
Empirical experiments conducted on the Hateful Meme Dataset Kiela et al. (2020) validate that HateSieve not only surpasses existing large multimodal models (LMMs) in performance—with fewer parameters—but also capable of interpreting and segmenting the visual and textual elements of multimodal memes to effectively identify discriminatory content.
@@ 2 Related Work
Existing safety filters for Large Language Models (LLMs) and LMMs typically comprise Alignment Ghafouri et al. (2023); Touvron et al. (2023); Rafailov et al. (2024); Wu et al. (2024), Inference Guidance Bai et al. (2022); Chiang et al. (2023); Zhang et al. (2023), and Input/Output Filter components Alon and Kamfonas (2023); Hu et al. (2023). Alignment involves fine-tuning LLMs to meet safety objectives using methods such as reinforcement learning from human feedback (RLHF) that optimize models based on safety data and human preferences. Inference guidance steers models towards generating safer responses through system prompts and token selection adjustments during generation. Input&Output filters detect and manage harmful content. However, these methods are primarily designed for unimodal content and struggle to adapt to multimodal content, such as confounder memes.
@@
Another category of methods uses representation learning and includes lightweight methods such as MOMENTA Pramanick et al. (2021), PromptHate Cao et al. (2022), and HateClipper Kumar and Nandakumar (2022b). MOMENTA constructs intra-modality attention by integrating external facial recognition data and background knowledge with the CLIP model. PromptHate converts images into text and then classifies them using a language model. HateClipper creates an image-text interaction matrix to fuse multimodal information. These methods enable straightforward classification with fewer parameters, but they offer limited interpretability of their classifications.
In contrast, our HateSieve framework generates context-aware image and text embeddings that enable effective meme segmentation and provide visual interpretation, while delivering classification performance comparable to existing methods.
@@ 3 Methodology
The HateSieve workflow involves: 1) Generating a triplet dataset with the Contrastive Meme Generator. 2) Pre-training the ITA module using the triplet dataset. 3) Extracting attention maps and performing segmentation with the pre-trained ITA. 4) Fine-tuning the ITA for hate speech classification.
Figure 3: The overall structure of our contrastive meme generator.
@@ 3.1 Contrastive Meme Generator
Our Contrastive Meme Generator is designed to produce both non-hateful and hateful versions of any given meme $\{(I_{i},M_{i},T_{i})\}_{i=1}^{N}$, where $I_{i}\in\mathbb{R}^{H\times W\times C}$ is the image pixels of the meme, $M_{i}\in\mathbb{R}^{H\times W}$ is the caption mask, and $T_{i}$ is the caption overlaid on the meme. These non-hateful and hateful versions are then used for subsequent contrastive learning. The first step in our Contrastive Meme Generator (CMG) is modality separation. By isolating the caption from the meme, we remove text borders and artifacts that may interfere with the image information, ensuring clean image content. Specifically, we employ the LaMA image Suvorov et al. (2021) inpainting pipeline to extract the pure image content $I_{i}^{\prime}=f_{\mathrm{LaMA}}(I_{i},M_{i})$ from the meme.
@@
Constructing a hateful version of a meme presents significant challenges due to the absence of direct annotations regarding ethnic groups, religious affiliations, social groups, or cultural identities in the meme $(I_{i},T_{i})$. This lack of explicit metadata complicates the generation of semantically similar hateful memes. To address this, we selected the largest available multimodal hate speech dataset, MMHS150k Gomez et al. (2020), focusing specifically on its “hateful” category for our reference dataset. We then employed the CLIP image encoder Radford et al. (2021) to derive embeddings for all image representations $\mathbf{I}^{-}\in\mathbb{R}^{N\times d}$, where $N$ represents the total number of memes in the dataset, and $d$ is the dimension of the embeddings. These embeddings serve as our reference embedding database. For each image $I_{i}$, we apply FAISS Douze et al. (2024) to find the most similar image based on Euclidean distance:
$k^{*}=\arg\min_{\mathbf{I}^{-}_{k}}\|f_{\mathrm{CLIP}}(I_{i})-\mathbf{I}^{-}_{k}\|_{2}$
where $f_{\mathrm{CLIP}}(I_{i})$ is the embedding of $I_{i}$. The closest negative pair, $(I^{-}_{k^{*}},T^{-}_{k^{*}})$, corresponds to the original pair $(I_{i},T_{i})$.
@@ 3.2 Triplet Dataset Generation
Our study constructs triplets of pairs for contrastive learning, each composed of an original meme $(I_{i},T_{i})$ and its two variations:
@@
To ensure the differentiation between hateful and non-hateful memes while preserving semantic coherence, each component within a meme—-either an image $I_{i}$ or text $T_{i}$-—is subjected to a pre-filtering process to identify potentially offensive or controversial content. Initially, each meme undergoes a dual filtering process as follows:
  • Text Filtering: Employing the IMSyPP Filter, the textual content $T_{i}$ is evaluated for offensive or violent/controversial material, yielding a classification $y_{i}^{T}$, where 1 indicates offensive (controversial) and 0 indicates non-offensive (safe).
  • Image Filtering: Using the NSFW filter from Stable Diffusion, images $I_{i}$ are assessed for inappropriate content like nudity or violence, with $y_{i}^{I}$ being 1 for NSFW (controversial) content and 0 for safe content.
Figure 4: The overall structure of our triplet dataset.
As shown in Figure 4, based on the filtering results, we construct the triplet dataset with the following process. For non-hateful pairs, designated as $(I_{i}^{\text{Non-Hate}},T_{i}^{\text{Non-Hate}})$, we sample from several combinations:
  • $(I_{i}^{+},T_{i}^{+})$ comprises both image and text generated by the CMG, ensuring that both elements are safe.
  • $(I_{i},T_{i}^{+})$ includes the original image deemed safe by the NSFW filter, paired with text that has been generated by the CMG as safe text.
  • $(I_{i}^{+},T_{i})$ features an image generated by the CMG, paired with the original text that has been classified as safe.
Conversely, for hateful pairs $(I_{i}^{\text{Hate}},T_{i}^{\text{Hate}})$, we sample from the following scenarios:
  • $(I_{i}^{-},T_{i}^{-})$ where both the image and text are generated by the CMG as controversial, embodying content flagged for potential offensiveness.
  • $(I_{i},T_{i}^{-})$ utilizes the original image classified as controversial, paired with controversially generated text by the CMG.
  • $(I_{i}^{-},T_{i})$ combines a controversially generated image by the CMG with the original text classified as controversial.
Figure 5: An overview of the ITA module: The representation $\mathbf{H}_{i}$ is produced for contrastive learning pre-training and subsequent classification tasks. Additionally, the attention map from the self-attention layer is utilized for segmenting hate components.
@@ 3.3 Image-Text Alignment Module
For each meme $(I_{i},T_{i})$, our ITA module is designed to derive a token-level, context-aware representation that integrates both the image and the text components, as illustrated in Figure 5. The process unfolds as follows:
@@
$\mathbf{H}_{i}=f_{\mathrm{DC}}\left(\mathrm{LN}\left([\mathbf{I}^{\mathrm{Align}}_{i},\mathbf{T}^{\mathrm{Align}}_{i}]\oplus[\mathbf{I}^{\mathrm{Pool}}_{i},\mathbf{T}^{\mathrm{Pool}}_{i}]\right)\right)$ (3)
where $\oplus$ denotes the operation for residual connection, $\mathrm{LN}$ represents layer normalization, and $f_{\mathrm{DC}}$ is the decoder, which is constructed with a multi-layer perception (MLP).
@@ 3.5 Hate Component Segmentation
Our hate component segmentation is structured as follows: After the ITA module is pre-trained via contrastive learning, it can process any given meme $(I_{i},T_{i})$ to extract a series of self-attention maps $\{\mathrm{Attn}_{i}\}_{\ell=1}^{L}$ from all layers. We begin by averaging these self-attention maps across layers to obtain $\mathrm{Attn}_{i}^{\prime}$. We then isolate the image attention map $\mathrm{Attn}_{l_{j},l_{t}}^{\prime}$ and the text attention map $\mathrm{Attn}_{l_{t},l_{j}}^{\prime}$, where $11<l_{j}<L_{I}+$ 1222We begin with the second image representation because the first one is a class embedding configured in CLIP, which is not applicable for segmentation purposes. and $L_{I}+1<l_{t}<L_{T}$. Here, $l_{j}$ denotes the $j$-th image patch among a total of $L_{I}$ patches, and $l_{t}$ indicates the $t$-th text token within a maximum of $L_{T}$ text tokens.
@@
Figure 6: Hate component segmentation process with attention map.
To construct an image segmentation map, we employ bilinear interpolation to upscale the $L_{I}\times L_{I}$ patch-level attention maps to $H\times W$pixel-level resolution, facilitating detailed visual analysis of the meme components. As for the text segmentation, we select the Top-$k$ tokens based on the attention scores per token, which allows for precise identification and analysis of the most contextually significant textual elements within the meme.

### v2

# A Context-Aware Contrastive Learning Framework for Hateful Meme Detection and Segmentation

@@ Abstract
Amidst the rise of Large Multimodal Models (LMMs) and their widespread application in generating and interpreting complex content, the risk of propagating biased and harmful memes remains significant. Current safety measures often fail to detect subtly integrated hateful content within “Confounder Memes”. To address this, we introduce HateSieve, a new framework designed to enhance the detection and segmentation of hateful elements in memes. HateSieve features a novel Contrastive Meme Generator that creates semantically correlated memes, a customized triplet dataset for contrastive learning, and an Image-Text Alignment module that produces context-aware embeddings for accurate meme segmentation. Empirical experiments show that HateSieve not only surpasses existing LMMs in performance with fewer trainable parameters but also offers a robust mechanism for precisely identifying and isolating hateful content. Caution: Contains academic discussions of hate speech; viewer discretion advised.
@@ 1 Introduction
The emergence of large multimodal models (LMMs), such as GPT-4V Achiam et al. (2023), Stable Diffusion Rombach et al. (2022), and DALL·E Ramesh et al. (2022), has ushered in a new era in which people increasingly rely on these models to generate and interpret visual and textual information. While these services simplify access to information—as illustrated in Figure 1—they also introduce risks of unregulated content that could distort public perception and harm social groups Su et al. (2023); Qu et al. (2023); Chin et al. (2023); Qu et al. (2024); Meng Huat Tiong et al. (2024); Lin et al. (2024). To address this risk, current LMM platforms implement safety filters, incorporating Alignment Ghafouri et al. (2023), Inference Guidance Chiang et al. (2023), and Input&Output Filter Alon and Kamfonas (2023) to detect and eliminate offensive or inappropriate components in both images and text.
However, these safety filters face challenges in identifying “Confounder Memes” Kiela et al. (2020); Mei et al. (2023), which deliberately combine visual and textual elements to convey biased and discriminatory messages. These memes may lack overtly offensive content in their individual components but deliver harmful messages through their combined presentation, making them difficult to detect using conventional safety mechanisms.
@@
Figure 2: An overview of our HateSieve framework. In Phase 1, we use the Contrastive Meme Generator to create a context-correlated triplet meme dataset, pre-training the model via contrastive learning. In Phase 2, we add a classification head and fine-tune the model on the downstream task, enabling it to classify memes while producing segmentation maps of hateful content.
A straightforward solution involves supervised fine-tuning (SFT) of LMMs Lin et al. (2024) to recognize hateful semantics in confounder memes. Nevertheless, this approach encounters several obstacles: (1) the scarcity of pairwise annotations in existing hateful meme datasets makes it difficult for models to accurately distinguish between hateful and non-hateful memes, especially when the differences are subtle; (2) deploying LMMs as safety filters alongside their regular online service usage111“Online service” refers to real-time applications like chatbots, virtual assistants, and image recognition platforms that use LMMs. is computationally intensive and non-trivial Lin et al. (2024). Alternatively, a lightweight classifier Kumar and Nandakumar (2022a); Mei et al. (2023) could be trained from scratch using a specialized hateful meme dataset, but this method suffers from limited interpretability and cannot provide detailed segmentation to explain its classifications.
To address these challenges, we introduce HateSieve, a novel framework for detecting hateful memes, as detailed in Figure 2. HateSieve mitigates the scarcity of detailed annotations by incorporating a Contrastive Meme Generator (CMGen), which constructs contextually correlated triplet datasets from existing memes. CMGen generates semantically similar but contrasting hateful and non-hateful memes within the same contextual scenarios, enabling the model to implicitly learn the subtle differences between hateful and non-hateful content. To facilitate detailed meme segmentation, HateSieve incorporates an Image-Text Alignment (ITA) module coupled with a frozen CLIP model. By pre-training on CMGen-generated triplets using contrastive learning in Phase 1, the ITA module develops context-aware attention maps that effectively segment both image and text hateful elements within memes. In Phase 2, the ITA module incorporates a fine-tuned classification head, leveraging its learned representations for hateful content classification. Empirical experiments conducted on various datasets validate that HateSieve not only outperforms existing LMMs with fewer parameters but also excels in interpreting and segmenting the visual and textual components of multimodal memes to effectively identify discriminatory content. Our contributions are summarized as follows:
  • We introduce CMGen, which generates context-correlated triplet pairs, filling the gap where specific pairwise annotations are absent in existing hateful meme datasets.
  • We present the ITA module that efficiently produces context-aware attention maps for both images and texts. These maps significantly enhance the model’s ability to segment and identify discriminatory elements within memes.
@@ 2 Related Work
Existing safety filters for Large Language Models (LLMs) and LMMs typically comprise Alignment Ghafouri et al. (2023); Touvron et al. (2023); Rafailov et al. (2024); Wu et al. (2024), Inference Guidance Bai et al. (2022); Chiang et al. (2023); Zhang et al. (2023), and Input&Output Filter components Alon and Kamfonas (2023); Hu et al. (2023). Alignment involves fine-tuning LLMs to meet safety objectives using methods such as reinforcement learning from human feedback (RLHF) that optimize models based on safety data and human preferences. Inference guidance steers models towards generating safer responses through system prompts and token selection adjustments during generation. Input&Output filters detect and manage harmful content. However, these methods are primarily designed for unimodal content and struggle to adapt to multimodal content, such as confounder memes.
Figure 3: Structure of CMGen: From any meme—including image $I_{i}$, text $T_{i}$, and caption mask $M_{i}$—CMGen generates corresponding hateful and non-hateful counterparts.
@@
Another category of methods uses representation learning and includes lightweight methods such as MOMENTA Pramanick et al. (2021), PromptHate Cao et al. (2022), and HateClipper Kumar and Nandakumar (2022b). MOMENTA constructs intra-modality attention by integrating external facial recognition data and background knowledge with the CLIP model. PromptHate converts images into text and then classifies them using a language model. HateClipper creates an image-text interaction matrix to fuse multimodal information. These methods enable straightforward classification with fewer parameters, but they offer limited interpretability of their classifications. In contrast, our HateSieve framework generates context-aware attention map that enable effective meme segmentation and provide visual interpretation, while delivering classification performance comparable to existing methods.
Figure 4: An overview of the triplet dataset generation process and our Image-Text Alignment (ITA) module.
@@ 3 Methodology
The HateSieve workflow involves: 1) Generating a triplet dataset with the CMGen. 2) Pre-training the ITA module using the triplet dataset. 3) Extracting attention maps and performing segmentation with the pre-trained ITA. 4) Fine-tuning classification head for hateful content classification.
@@ 3.1 Contrastive Meme Generator
As shown in Figure 3, our CMGen is designed to produce both non-hateful and hateful versions of any given meme $\{(I_{i},M_{i},T_{i})\}_{i=1}^{N}$, where $I_{i}\in\mathbb{R}^{H\times W\times C}$ is the image pixels of the meme, $M_{i}\in\mathbb{R}^{H\times W}$ is the caption mask, and $T_{i}$ is the caption overlaid on the meme. These non-hateful and hateful versions are then used for subsequent contrastive learning. The first step in our CMGen is modality separation. By isolating the caption from the meme, we remove text borders and artifacts that may interfere with the image information, ensuring clean image content. Specifically, we apply the LaMA image Suvorov et al. (2021) inpainting pipeline to extract the pure image content $I_{i}^{\prime}=f_{\mathrm{LaMA}}(I_{i},M_{i})$ from the meme.
@@
Constructing a hateful version of a meme $(I_{i}^{-},T_{i}^{-})$ presents significant challenges due to the absence of direct annotations regarding ethnic groups, religious affiliations, social groups, or cultural identities in the original meme $(I_{i},T_{i})$. This lack of explicit metadata complicates the generation of semantically similar hateful memes. To address this issue, we selected the largest available multimodal hate speech dataset, MMHS150k Gomez et al. (2020), focusing specifically on its “hateful” category to serve as our reference dataset $\mathcal{R}={(I_{r}^{-},T_{r}^{-})}_{r=1}^{N_{\mathcal{R}}}$, where $N_{\mathcal{R}}$ denotes the number of memes in the reference dataset.
For each purified image $I_{i}^{\prime}$ of $I_{i}$, we aim to find the most similar hateful image [^2] We only use image embeddings for similarity search because the text in memes often lacks explicit social group features, making it less effective for finding semantically similar hateful pairs. from the reference dataset $\mathcal{R}$. We utilize the CLIP image encoder Radford et al. (2021) $f_{\mathrm{CLIP}}$ to compute the embeddings of both the purified image and the images in the reference dataset. Using FAISS Douze et al. (2024) for efficient similarity search, we find the index $r^{*}$ of the most similar image based on Euclidean distance:
$r^{*}=\arg\min_{r\in\{1,\dots,N_{R}\}}\left\|f_{\mathrm{CLIP}}(I_{i}^{\prime})-f_{\mathrm{CLIP}}(I_{r}^{-})\right\|_{2}$
The closest hateful pair $(I_{r^{*}}^{-},T_{r^{*}}^{-})$ from the reference dataset is then used as the hateful version of our original meme.
@@ 3.2 Triplet Dataset Generation
Our study constructs triplets of meme pairs for contrastive learning, each composed of an original meme $(I_{i},T_{i})$ and its two variations:
@@
To distinguish between hateful and non-hateful content while maintaining semantic coherence, each meme component—the image $I_{i}$ and the text $T_{i}$—undergoes a pre-filtering process to identify potentially offensive or controversial material. Specifically, each meme is filtered as follows:
  • Text Filtering: Using the IMSyPP Filter Kralj Novak et al. (2022), we evaluate the text $T_{i}$ for offensive or controversial content, assigning a label $y_{i}^{T}$, where $y_{i}^{T}=1$ indicates offensive content and $y_{i}^{T}=0$ indicates non-offensive content.
  • Image Filtering: Employing the NSFW filter from Stable Diffusion, we assess the image $I_{i}$ for inappropriate content such as nudity or violence, resulting in a label $y_{i}^{I}$, where $y_{i}^{I}=1$ denotes NSFW content and $y_{i}^{I}=0$ denotes safe content.
As illustrated in Figure 4, we construct the triplet dataset based on these filtering results:
Non-Hateful Pairs $(I_{i}^{\text{Non-Hate}},T_{i}^{\text{Non-Hate}})$:
We sample from the following combinations to ensure both image and text are non-offensive:
  • $(I_{i}^{+},T_{i}^{+})$: The non-hateful meme generated by CMGen without any offensive contents.
  • $(I_{i},T_{i}^{+})$: The original image ($y_{i}^{I}=0$) is paired with safe text generated by CMGen.
  • $(I_{i}^{+},T_{i})$: A safe image generated by CMGen is paired with the original text ($y_{i}^{T}=0$).
Hateful Pairs $(I_{i}^{\text{Hate}},T_{i}^{\text{Hate}})$:
We sample from the following combinations to include offensive elements as hateful meme:
  • $(I_{i}^{-},T_{i}^{-})$: The hateful meme generated by CMGen that contains offensive content.
  • $(I_{i},T_{i}^{-})$: The original image ($y_{i}^{I}=1$) is paired with offensive text from CMGen.
  • $(I_{i}^{-},T_{i})$: An offensive image generated by CMGen is combined with the original text that contains offensive content ($y_{i}^{T}=1$).
@@ 3.3 Image-Text Alignment Module
For each meme $(I_{i},T_{i})$, our ITA module is designed to derive a token/patch-level, context-aware representation that integrates both the image and the text components, as illustrated in Figure 4. The process unfolds as follows:
@@
$\mathbf{H}_{i}=f_{\mathrm{Decoder}}\left([\mathbf{I}^{\mathrm{Align}}_{i},\mathbf{T}^{\mathrm{Align}}_{i}]\oplus[\mathbf{I}^{\mathrm{Pool}}_{i},\mathbf{T}^{\mathrm{Pool}}_{i}]\right)$ (3)
where $\oplus$ denotes the operation for residual connection and $f_{\mathrm{Decoder}}$ denotes the decoder, which incorporates a Multilayer Perceptron (MLP) module for dimensionality reduction.
@@ 3.5 Hate Component Segmentation
Our hate component segmentation is structured as follows: After the ITA module is pre-trained via contrastive learning, it can process any given meme $(I_{i},T_{i})$ to extract a series of self-attention maps $\{\mathrm{Attn}_{i}\}_{\ell=1}^{L}$ from all layers. We begin by averaging these self-attention maps across layers to obtain $\mathrm{Attn}_{i}^{\prime}$. We then isolate the image attention map $\mathrm{Attn}_{l_{j},l_{t}}^{\prime}$ and the text attention map [^3] We begin with the second image representation because the first one is a class embedding configured in CLIP, which is not applicable for segmentation purposes. $\mathrm{Attn}_{l_{t},l_{j}}^{\prime}$, where $11<l_{j}<L_{I}+$ 1 and $L_{I}+1<l_{t}<L_{T}$. Here, $l_{j}$ denotes the $j$-th image patch among a total of $L_{I}$ patches, and $l_{t}$ indicates the $t$-th text token within a maximum of $L_{T}$ text tokens.
@@
To construct an image segmentation map, we employ bilinear interpolation to upscale the $L_{I}\times L_{I}$ patch-level attention maps to $H\times W$pixel-level resolution, facilitating detailed visual analysis of the meme components. As for the text segmentation, we select the Top-$k$ tokens based on the attention scores per token, which allows for precise identification and analysis of the most contextually significant textual elements within the meme. Details of the segmentation process are in Appendix A.1.